A tip jar at a diner contained $10.00 in coins. It contained 10 more dimes than quarters and twice as many nickels as quarters. How many dimes were in the tip jar?
Answer by ikleyn(52786)

.
A tip jar at a diner contained $10.00 in coins. It contained 10 more dimes than quarters
and twice as many nickels as quarters. How many dimes were in the tip jar?
~~~~~~~~~~~~~~~~~~~~~~~


d  dimes,

(d-10)  quarters,

2*(d-10)  nickels.


The total money equation


    5*2(d-10) + 10d + 25*(d-10) = 1000  cents.


Simplify and find d


    10(d-10) + 10d + 25(d-10)   = 1000

    10d - 100 + 10d + 25d - 250 = 1000

            45d                 = 1000 + 100 + 250 

            45d                 = 1350

              d                 = 1350/45 = 30.


ANSWER.  30 dimes.

Solved.